Solve the following equations. 
14 y – 8 = 13
Solve the following equations. 
14 y – 8 = 13

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

14y–8=13
14y = 13+8
14y=21
y = 21/14
y=3/2

Thanks
Shraddha Singh
2 years ago
Dear Student,
We can solve it as follows :
 
14 y - 8 = 13
14 y = 13 + 8
14 y = 21
Y = 21/14
Y = 3/2
 
Hope it helps!
Thank you ✌️